# ReQISC: A Reconfigurable Quantum Computer Microarchitecture and Compiler Co-Design

### v1

@@ Appendix A Proof of the Algorithm 1 (genAshN gate scheme)
Takeaways Since the presented scheme could also be regarded as a “generalized” version of the AshN scheme and their proof and solving thoughts are sort of similar, we also call the ReQISC microarchitecture as “genAshN” scheme throughout this appendix.
Figure 15. Overview of the proof.
@@ A.1.1. Hamiltonian Canonicalization
$\small H[a,b,c]:=a\cdot\sigma_{X}\otimes\sigma_{X}+b\cdot\sigma_{Y}\otimes\sigma_{Y}+c\cdot\sigma_{Z}\otimes\sigma_{Z}.$
We follow the canonicalization treatment in  (dur2001entanglement) and only consider two-qubit Hamiltonians of the form $H[a,b,c]$ where $a\geq b\geq|c|,a>0$. Furthermore, we can rescale any Hamiltonian as long as the corresponding gate time is rescaled accordingly. We thus identify canonical Hamiltonian coefficients $H[a,b,c]\sim H[ka,kb,kc]$ for all $k>0$.
@@ A.1.2. Two-qubit gate Canonicalization
Through the well-known canonical decomposition (or KAK decomposition) (zhang2003geometric), any two-qubit unitary can be determined by a unique Weyl coordinate $(x,y,z)$ such that
$U=g\cdot(V_{1}\otimes V_{2})e^{i(x\,XX+y\,YY+z\,ZZ)}(V_{3}\otimes V_{4}),\,g\in\{1,i\},\,\pi/4\geq x\geq y\geq|z|$
The canonical coordinate $(x,y,z)$ which determines a class of two-qubit gates locally equivalent to the so-defined canonical gate $\mathrm{Can}(x,y,z)=e^{i\vec{\eta}\cdot\vec{\Sigma}}=e^{i(x\,XX+y\,YY+z\,ZZ)}$ is confined to a tetrahedron dubbed the Weyl chamber [^2] Note that for the convenience of proof, herein we select the definition of Canonical gate $(x,y,z)\sim e^{i\vec{\eta}\cdot\vec{\Sigma}}$ that is different from the one $(x,y,z)\sim e^{-i\vec{\eta}\cdot\vec{\Sigma}}$ in the our paper submission..
@@ A.1.3. Problem Statement
Given $H[a,b,c]$, the minimum evolution time $\tau_{\text{opt}}$ (hammerer2002characterization) to achieve the Weyl chamber coordinate $(x,y,z)$, which indicates a class of two-qubit gates locally equivalent to the so-defined canonical gate $\mathrm{Can}(x,y,z)=e^{-i(x\,XX+y\,YY+z\,ZZ)}$ with $\pi/4\geq x\geq y\geq|z|$, by alternately applying $H[a,b,c]$ and single-qubit Hamiltonians $H_{0}\otimes H_{1}$ satisfies $\tau_{\text{opt}}=\min\{\tau_{1},\tau_{2}\}$, where
@@
$\small W_{ext}:=\{(x,y,z)\in\mathbb{R}^{3}|0\leq x\leq\pi/2,0\leq y\leq\min\{x,\pi/2-x\},-y\leq z\leq y\}.$
@@
$A_{\tau}:=\left(ND(a,b,c;\tau)\cup EA_{+}(a,b,c;\tau)\cup EA_{-}(a,b,c;\tau)\right),$
where
$ND(a,b,c;\tau)$ $:=\left\{(x,y,z)|x=a\tau,y\pm z\in[0,(b\pm c)\tau]\right\},$
$EA_{+}(a,b,c;\tau)$ $:=\left\{(x,y,z)|x+y+z=(a+b+c)\tau,a\tau\geq x\geq y\geq|z|,z\geq c\tau\right\},$
$EA_{-}(a,b,c;\tau)$ $:=\left\{(x,y,z)|x+y-z=(a+b-c)\tau,a\tau\geq x\geq y\geq|z|,z\leq c\tau\right\}.$
@@
$\tau=\min\left\{t\mid(x,y,z)\in\bigcup_{t^{\prime}=0}^{t}A_{t}\right\}.$
It can be verified that this is equivalent to $\tau$ being the minimum number that simultaneously satisfies $a\tau\geq x,(a+b+c)\tau\geq x+y+z$ and $(a+b-c)\tau\geq x+y-z$, for any coordinate $0\leq|z|\leq y\leq x$. This helps us formulate the optimal gate time as follows:
@@
$\tau_{\text{opt}}(a,b,c;x,y,z):=\min\left\{\max\left\{\frac{x}{a},\frac{x+y+z}{a+b+c},\frac{x+y-z}{a+b-c}\right\},\max\left\{\frac{\frac{\pi}{2}-x}{a},\frac{\frac{\pi}{2}-x+y-z}{a+b+c},\frac{\frac{\pi}{2}-x+y+z}{a+b-c}\right\}\right\}$
@@
$\exp\{-i\tau_{\text{opt}}(a,b,c;x,y,z)(H[a,b,c]+(\Omega_{1}+\Omega_{2})X\otimes I+(\Omega_{1}-\Omega_{2})I\otimes X+\delta(Z\otimes I+I\otimes Z))\}$
@@
For any canonical Hamiltonian coefficients $a\geq b\geq|c|$, any $\tau\in(0,\frac{\pi}{2a}]$, for any Weyl chamber coordinates $(x,y,z)\in ND(a,b,c;\tau)\cap W_{ext}$, there exist $r_{1},r_{2}\geq 0$ such that the interaction coefficients for $U:=\exp\{-i\cdot(H[a,b,c]+r_{1}\sigma_{X}\otimes I+r_{2}I\otimes\sigma_{X})\cdot\tau\}$ are $(x,y,z)$.
@@
For any canonical Hamiltonian coefficients $a\geq b\geq|c|$, $\tau\geq 0$, for any Weyl chamber coordinates $(x,y,z)\in EA_{+}(a,b,c;\tau)\cap W_{ext}$, there exist $\Omega,\delta\geq 0$ such that the interaction coefficients for $U:=\exp\{-i\cdot(H[a,b,c]+\Omega(\sigma_{X}\otimes I+I\otimes\sigma_{X})+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\cdot\tau\}$ are $(x,y,z)$.
@@
For any canonical Hamiltonian coefficients $a\geq b\geq|c|$, $\tau\geq 0$, for any Weyl chamber coordinates $(x,y,z)\in EA_{-}(a,b,c;\tau)\cap W_{ext}$, there exist $\Omega,\delta\geq 0$ such that the interaction coefficients for $U:=\exp\{-i\cdot(H[a,b,c]+\Omega(\sigma_{X}\otimes I-I\otimes\sigma_{X})+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\cdot\tau\}$ are $(x,y,z)$.
@@
For every $(x,y,z)\in EA_{-}(a,b,c;\tau)$, we want to prove that there exist $\Omega,\delta>0$ such that the interaction coefficients for $U:=\exp\{-i\cdot(H[a,b,c]+\Omega(\sigma_{X}\otimes I-I\otimes\sigma_{X})+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\cdot\tau\}$ are $(x,y,z)$. We consider a similar dual problem of generating the Weyl chamber coordinates (x,y,z)(x,y,-z) with single-qubit time-invariant Hamiltonians using H[a,b,c]H[a,b,-c]. It can be verified that $(x,y,-z)\in EA_{+}(a,b,-c;\tau)$; therefore there exist $\Omega^{\prime}$ and $\delta^{\prime}$ such that
$\small U^{\prime}:=\exp\{-i\cdot(H[a,b,-c]+\Omega^{\prime}(\sigma_{X}\otimes I+I\otimes\sigma_{X})+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\cdot\tau\}$
@@
$(\sigma_{Z}\otimes I)U^{\prime}(\sigma_{Z}\otimes I)$ $=\exp\{-i\cdot(H[-a,-b,-c]+\Omega^{\prime}(-\sigma_{X}\otimes I+I\otimes\sigma_{X})+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\cdot t\}$
$=\exp\{-i\cdot(H[a,b,c]+\Omega^{\prime}(\sigma_{X}\otimes I-I\otimes\sigma_{X})-\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\cdot\tau\}^{{\dagger}}$
@@
$\small\exp\{-i\cdot(H[a,b,c]+\Omega^{\prime}(\sigma_{X}\otimes I-I\otimes\sigma_{X})-\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\cdot\tau\}$
@@ A.1.4. Transforming Weyl chamber coordinates to spectrum
$\sim\exp\{i\vec{\eta}\cdot\vec{\Sigma}\}\cdot\sigma_{Y}\otimes\sigma_{Y},$
@@
$\small\mathcal{M}:=\frac{1}{\sqrt{2}}\begin{bmatrix}1&0&0&i\\ 0&i&1&0\\ 0&i&-1&0\\ 1&0&0&-i\end{bmatrix}$.
@@
$\small(\mathcal{M}^{\dagger}U\mathcal{M})^{T}=\mathcal{M}^{T}U\bar{\mathcal{M}}.$
@@
$\small D^{\prime}=\begin{bmatrix}1&&&\\ &-1&&\\ &&1&\\ &&&-1\end{bmatrix}$.
@@
$\small\mathcal{M}^{\dagger}U\mathcal{M}D^{\prime}=ODO^{T}\sim D.$
@@
$:=$ $\exp\{-i\tau(H[a,b,c]+\Omega_{1}(\sigma_{X}\otimes I+I\otimes\sigma_{X})+\Omega_{2}(\sigma_{X}\otimes I-I\otimes\sigma_{X})+\delta(\omega_{Z}\otimes I+I\otimes\omega_{Z}))\}\cdot(\sigma_{Y}\otimes\sigma_{Y}).$
@@ A.2. Canonicalization for genAshN-ND
$H_{ND}=$ $(H\otimes H)(H[a,b,c]+\Omega_{1}(\sigma_{X}\otimes I+I\otimes\sigma_{X})+\Omega_{2}(\sigma_{X}\otimes I-I\otimes\sigma_{X}))(H\otimes H)$
$=$ $\begin{pmatrix}a+2\Omega_{1}&&&c-b\\ &-a+2\Omega_{2}&b+c&\\ &b+c&-a-2\Omega_{2}&\\ c-b&&&a-2\Omega_{1}\end{pmatrix}$.
@@
$=$ $\exp\{-iH_{ND}\tau\}(\sigma_{Y}\otimes\sigma_{Y})$
$=$ $\begin{pmatrix}\frac{ie^{-ia\tau}(c-b)\sin S_{1}\tau}{S_{1}}&&&\frac{e^{-ia\tau}(-S_{1}\cos S_{1}\tau+i\Omega_{1}\sin S_{1}\tau)}{S_{1}}\\ &\frac{-ie^{ia\tau}(b+c)\sin S_{2}\tau}{S_{2}}&\frac{e^{ia\tau}(S_{2}\cos S_{2}\tau-i\Omega_{2}\sin S_{2}\tau)}{S_{2}}&\\ &\frac{e^{ia\tau}(S_{2}\cos(S_{2}\tau)+i\Omega_{2}\sin S_{2}\tau)}{S_{2}}&\frac{-ie^{ia\tau}(b+c)\sin S_{2}\tau}{S_{2}}&\\ \frac{e^{-ia\tau}(-S_{1}\cos S_{1}\tau-i\Omega_{1}\sin S_{1}\tau)}{S_{1}}&&&\frac{ie^{-ia\tau}(c-b)\sin S_{1}\tau}{S_{1}}\end{pmatrix}$
where $S_{1}:=\sqrt{4\Omega_{1}^{2}+(b-c)^{2}},S_{2}:=\sqrt{4\Omega_{2}^{2}+(b+c)^{2}}$. The spectrum of $V$ is then $\{ie^{-i(at\pm\theta_{1})},-ie^{i(at\pm\theta_{2})},\}$, where $\cos\theta_{1}:=\frac{(c-b)\sin(S_{1}t)}{S_{1}}$ and $\cos\theta_{2}:=\frac{(b+c)\sin S_{2}t}{S_{2}}$. It can be verified that such a gate corresponds to Weyl chamber coordinates $(at,y,z)$, where
$\small\sin(y+z)=\frac{(b+c)\sin S_{2}t}{S_{2}},\sin(y-z)=\frac{(b-c)\sin S_{1}t}{S_{1}}.$

### v2

@@ Appendix A Proof of the Algorithm 1 (genAshN gate scheme)
Takeaways Since the presented scheme could also be regarded as a “generalized” version of the AshN scheme and their proof and solving thoughts are sort of similar, we also call the ReQISC microarchitecture as “genAshN” scheme throughout this appendix. Source code implementation, test files, and usage examples are available in the Zenodo archive https://zenodo.org/records/18163249.
@@ A.1.1. Hamiltonian Canonicalization
(8) $H[a,b,c]:=a\cdot\sigma_{X}\otimes\sigma_{X}+b\cdot\sigma_{Y}\otimes\sigma_{Y}+c\cdot\sigma_{Z}\otimes\sigma_{Z}.$
We follow the canonicalization treatment in  (Dür et al., 2001) and only consider two-qubit Hamiltonians of the form $H[a,b,c]$ where $a\geq b\geq|c|,\,a>0$. Furthermore, we can rescale any Hamiltonian as long as the corresponding gate time is rescaled accordingly. We thus identify canonical Hamiltonian coefficients $H[a,b,c]\sim H[ka,kb,kc]$ for all $k>0$.
@@ A.1.2. Two-qubit gate Canonicalization
Through the well-known canonical decomposition (or KAK decomposition) (Zhang et al., 2003), any two-qubit unitary can be determined by a unique Weyl coordinate $(x,y,z)$ such that
$U$ $=g\cdot(V_{1}\otimes V_{2})e^{i(x\,XX+y\,YY+z\,ZZ)}(V_{3}\otimes V_{4}),$
(9) $\quad g\in\{1,i\},\,\pi/4\geq x\geq y\geq|z|$
The canonical coordinate $(x,y,z)$ which determines a class of two-qubit gates locally equivalent to the so-defined canonical gate
(10) $\mathrm{Can}(x,y,z)=e^{i\vec{\eta}\cdot\vec{\Sigma}}=e^{i(x\,XX+y\,YY+z\,ZZ)}$
is confined to a tetrahedron dubbed the Weyl chamber [^1] Note that for the convenience of proof, herein we select the definition of Canonical gate $(x,y,z)\sim e^{i\vec{\eta}\cdot\vec{\Sigma}}$ that is different from the one $(x,y,z)\sim e^{-i\vec{\eta}\cdot\vec{\Sigma}}$ in the our paper submission..
@@ A.1.3. Problem Statement
Given $H[a,b,c]$, the minimum evolution time $\tau_{\text{opt}}$ (Hammerer et al., 2002) to achieve the Weyl chamber coordinate $(x,y,z)$, which indicates a class of two-qubit gates locally equivalent to the canonical gate $\mathrm{Can}(x,y,z)$, by alternately applying $H[a,b,c]$ and single-qubit Hamiltonians $H_{0}\otimes H_{1}$ satisfies $\tau_{\text{opt}}=\min\{\tau_{1},\tau_{2}\}$, where
@@
$W_{ext}:=\big\{$ $(x,y,z)\in\mathbb{R}^{3}\mid 0\leq x\leq\pi/2,$
(13) $0\leq y\leq\min\{x,\pi/2-x\},-y\leq z\leq y\big\}.$
Figure 17. Overview of the proof.
@@
(14) $A_{\tau}:=ND\cup EA_{+}\cup EA_{-},$
where we abbreviate $ND:=ND(a,b,c;\tau)$, $EA_{\pm}:=EA_{\pm}(a,b,c;\tau)$, and
(15) $ND$ $:=\{(x,y,z)\mid x=a\tau,\,y\pm z\in[0,(b\pm c)\tau]\},$
$EA_{+}$ $:=\{(x,y,z)\mid x+y+z=(a+b+c)\tau,$
(16) $\quad a\tau\geq x\geq y\geq|z|,\,z\geq c\tau\},$
$EA_{-}$ $:=\{(x,y,z)\mid x+y-z=(a+b-c)\tau,$
(17) $\quad a\tau\geq x\geq y\geq|z|,\,z\leq c\tau\}.$
@@
(18) $\tau=\min\left\{t\mid(x,y,z)\in\bigcup_{t^{\prime}=0}^{t}A_{t}\right\}.$
It can be verified that this is equivalent to $\tau$ being the minimum number that simultaneously satisfies
(19) $\begin{cases}a\tau\geq x,\\ (a+b+c)\tau\geq x+y+z,\\ (a+b-c)\tau\geq x+y-z,\end{cases}$
for any coordinate $0\leq|z|\leq y\leq x$. This helps us formulate the optimal gate time as follows:
@@
$\tau_{\text{opt}}$ $:=\min\{\tau_{1},\tau_{2}\},\text{ where}$
(20) $\tau_{1}$ $:=\max\left\{\frac{x}{a},\frac{x+y+z}{a+b+c},\frac{x+y-z}{a+b-c}\right\},$
$\tau_{2}$ $:=\max\left\{\frac{\frac{\pi}{2}-x}{a},\frac{\frac{\pi}{2}-x+y-z}{a+b+c},\frac{\frac{\pi}{2}-x+y+z}{a+b-c}\right\}$
@@
$\exp\{-i\tau_{\text{opt}}(H[a,b,c]+(\Omega_{1}+\Omega_{2})X\otimes I$
(21) $\quad+(\Omega_{1}-\Omega_{2})I\otimes X+\delta(Z\otimes I+I\otimes Z))\}$
@@
For any canonical Hamiltonian coefficients $a\geq b\geq|c|$, any $\tau\in(0,\frac{\pi}{2a}]$, for any Weyl chamber coordinates $(x,y,z)\in ND(a,b,c;\tau)\cap W_{ext}$, there exist $r_{1},r_{2}\geq 0$ such that the interaction coefficients for
(22) $U:=\exp\{-i\cdot(H[a,b,c]+r_{1}\sigma_{X}\otimes I+r_{2}I\otimes\sigma_{X})\cdot\tau\}$
are $(x,y,z)$.
@@
For any canonical Hamiltonian coefficients $a\geq b\geq|c|$, $\tau\geq 0$, for any Weyl chamber coordinates $(x,y,z)\in EA_{+}(a,b,c;\tau)\cap W_{ext}$, there exist $\Omega,\delta\geq 0$ such that the interaction coefficients for
$U:=\exp\big\{-i\cdot\big($ $H[a,b,c]+\Omega(\sigma_{X}\otimes I+I\otimes\sigma_{X})$
(23) $+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z})\big)\cdot\tau\big\}$
are $(x,y,z)$.
@@
For any canonical Hamiltonian coefficients $a\geq b\geq|c|$, $\tau\geq 0$, for any Weyl chamber coordinates $(x,y,z)\in EA_{-}(a,b,c;\tau)\cap W_{ext}$, there exist $\Omega,\delta\geq 0$ such that the interaction coefficients for
$U:=\exp\big\{-i\cdot\bigl($ $H[a,b,c]+\Omega(\sigma_{X}\otimes I-I\otimes\sigma_{X})$
(24) $+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z})\bigr)\cdot\tau\big\}$
are $(x,y,z)$.
@@
For every $(x,y,z)\in EA_{-}(a,b,c;\tau)$, we want to prove that there exist $\Omega,\delta>0$ such that the interaction coefficients for
$U:=\exp\big\{-i\cdot\big($ $H[a,b,c]+\Omega(\sigma_{X}\otimes I-I\otimes\sigma_{X})$
(25) $+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z})\big)\cdot\tau\big\}$
are $(x,y,z)$. We consider a similar dual problem of generating the Weyl chamber coordinates (x,y,z)(x,y,-z) with single-qubit time-invariant Hamiltonians using H[a,b,c]H[a,b,-c]. It can be verified that $(x,y,-z)\in EA_{+}(a,b,-c;\tau)$; therefore there exist $\Omega^{\prime}$ and $\delta^{\prime}$ such that
$U^{\prime}:=\exp\big\{-i\cdot\big($ $H[a,b,-c]+\Omega^{\prime}(\sigma_{X}\otimes I+I\otimes\sigma_{X})$
(26) $+\delta^{\prime}(\sigma_{Z}\otimes I+I\otimes\sigma_{Z})\big)\cdot\tau\big\}$
@@
$(\sigma_{Z}\otimes I)U^{\prime}(\sigma_{Z}\otimes I)$
$=\exp\{-i(H[-a,-b,-c]+\Omega^{\prime}(-\sigma_{X}\otimes I+I\otimes\sigma_{X})$
$\quad+\delta^{\prime}(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\tau\}$
$=\exp\{-i(H[a,b,c]+\Omega^{\prime}(\sigma_{X}\otimes I-I\otimes\sigma_{X})$
(27) $\quad-\delta^{\prime}(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\tau\}^{{\dagger}}$
@@
$\exp\big\{-i\cdot\big($ $H[a,b,c]+\Omega^{\prime}(\sigma_{X}\otimes I-I\otimes\sigma_{X})$
(28) $-\delta^{\prime}(\sigma_{Z}\otimes I+I\otimes\sigma_{Z})\big)\cdot\tau\big\}$
@@ A.1.4. Transforming Weyl chamber coordinates to spectrum
(29) $\sim\exp\{i\vec{\eta}\cdot\vec{\Sigma}\}\cdot\sigma_{Y}\otimes\sigma_{Y},$
@@
(30) $\mathcal{M}:=\frac{1}{\sqrt{2}}\begin{bmatrix}1&0&0&i\\ 0&i&1&0\\ 0&i&-1&0\\ 1&0&0&-i\end{bmatrix}$.
@@
(31) $(\mathcal{M}^{\dagger}U\mathcal{M})^{T}=\mathcal{M}^{T}U\bar{\mathcal{M}}.$
@@
(32) $D^{\prime}=\begin{bmatrix}1&&&\\ &-1&&\\ &&1&\\ &&&-1\end{bmatrix}$.
@@
(33) $\mathcal{M}^{\dagger}U\mathcal{M}D^{\prime}=ODO^{T}\sim D.$
@@
$:=\exp\{-i\tau(H[a,b,c]+\Omega_{1}(\sigma_{X}\otimes I+I\otimes\sigma_{X})$
$\quad+\Omega_{2}(\sigma_{X}\otimes I-I\otimes\sigma_{X})$
(34) $\quad+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}))\}\cdot(\sigma_{Y}\otimes\sigma_{Y}).$
@@ A.2. Canonicalization for genAshN-ND
$H_{ND}$ $=(H\otimes H)(H[a,b,c]+\Omega_{1}(\sigma_{X}\otimes I+I\otimes\sigma_{X})$
$\quad+\Omega_{2}(\sigma_{X}\otimes I-I\otimes\sigma_{X}))(H\otimes H)$
(35) $=\begin{pmatrix}a+2\Omega_{1}&&&c-b\\ &-a+2\Omega_{2}&b+c&\\ &b+c&-a-2\Omega_{2}&\\ c-b&&&a-2\Omega_{1}\end{pmatrix}$.
@@
(36) $=\exp\{-iH_{ND}\tau\}(\sigma_{Y}\otimes\sigma_{Y}).$
Define $S_{1}:=\sqrt{4\Omega_{1}^{2}+(b-c)^{2}}$, $S_{2}:=\sqrt{4\Omega_{2}^{2}+(b+c)^{2}}$. This can be written in block form (with $V_{ij}$ denoting entries):
$V_{11}$ $=\frac{ie^{-ia\tau}(c-b)\sin S_{1}\tau}{S_{1}},$
$V_{14}$ $=\frac{e^{-ia\tau}(-S_{1}\cos S_{1}\tau+i\Omega_{1}\sin S_{1}\tau)}{S_{1}},$
(37) $V_{22}$ $=\frac{-ie^{ia\tau}(b+c)\sin S_{2}\tau}{S_{2}},$
$V_{23}$ $=\frac{e^{ia\tau}(S_{2}\cos S_{2}\tau-i\Omega_{2}\sin S_{2}\tau)}{S_{2}},$
with $V_{41}=\bar{V}_{14}$, $V_{44}=V_{11}$, $V_{32}=\bar{V}_{23}$, $V_{33}=V_{22}$. The spectrum of $V$ is $\{ie^{-i(at\pm\theta_{1})},-ie^{i(at\pm\theta_{2})}\}$, where $\cos\theta_{1}:=\frac{(c-b)\sin(S_{1}t)}{S_{1}}$ and $\cos\theta_{2}:=\frac{(b+c)\sin S_{2}t}{S_{2}}$. It can be verified that such a gate corresponds to Weyl chamber coordinates $(at,y,z)$, where
$\sin(y+z)=\frac{(b+c)\sin S_{2}t}{S_{2}},\,\sin(y-z)=\frac{(b-c)\sin S_{1}t}{S_{1}}.$